# Language in the Flow of Time: Time-Series-Paired Texts Weaved into a Unified Temporal Narrative

### v2

@@ 5.1 Main Results
Time Series Forecasting. Table 2 presents the performance results for the time series forecasting task. For each dataset and time series model, we report the average performance across four different prediction lengths, and the full results for each prediction length are provided in Appendix E. For datasets with relatively few samples, such as Economy, we perform short-term forecasting with prediction lengths of $\{6,8,10,12\}$. In contrast, for datasets with a larger number of samples, such as Environment, we perform long-term forecasting with prediction lengths of $\{48,96,192,336\}$. From the results, compared to uni-modal modeling or the state-of-the-art baseline MM-TSFLib, our TaTS consistently achieves the best performance across all datasets. Notably, by plugging in TaTS to various existing time series models, it achieves an average performance improvement of over $5\%$ on 6 out of 9 datasets and delivers a remarkable performance boost of over $30\%$ on the Environment dataset, which contains the largest number of samples among all datasets. The results also demonstrate that TaTS is highly compatible with a wide range of existing time series forecasting models, including transformer-based models, linear models, and frequency-based models, consistently delivering performance improvements across all of them in both long-term forecasting and short-term forecasting tasks. We provide a visualization of the performance boost in Appendix E.2, showcasing the improvements achieved by different time series models on each dataset.
Table 3: Performance of time-series imputation task with concurrent texts. Full results are reported in Appendix E.4. Promotion: percentage reduction in MSE or MAE achieved by TaTS compared to the best-performing baseline.
@@
Time Series Imputation. We also evaluate the performance of our TaTS framework on the imputation task using the Climate, Economy, and Traffic datasets, each with an imputation length of 24. Since the original MM-TSFLib only supports forecasting tasks, we extend it by applying a similar linear interpolation approach to serve as a baseline. For this evaluation, we select one representative time series model from each category and present the results in Table 3. The results demonstrate that TaTS consistently enhances the imputation capabilities of existing time series models, achieving improvements of up to $30\%$ compared to baseline methods.
From the above results, it is evident that leveraging the text modality provides significant benefits over numerical-only unimodal modeling when paired texts are available. Our proposed TaTS framework effectively captures the temporal properties encoded in paired texts, achieving notable improvements over baseline methods that disregard the positional information inherent in the texts paired with time series.
Table 4: TT-Wasserstein measure of Time-MMD [35] datasets and their random shuffle.
@@
Ratio of Original to Shuffled TT-Wasserstein (%) 26.8 72.4 22.3 91.5 38.2 34.6 97.1 85.6 83.6
Avg. Improvement of TaTS over uni-modal (%) 20.70 18.03 64.80 4.05 10.99 16.78 3.62 19.51 36.00
Table 5: TaTS compared several baselines.
Dataset Climate Security Traffic
Methods MSE MAE MSE MAE MSE MAE
iTransformer [2024c] 1.127 0.843 113.57 5.698 0.203 0.228
N-BEATS [2020] 1.123 0.894 131.91 6.866 0.352 0.409
N-HiTS [2022] 1.114 0.881 135.41 6.896 0.295 0.372
TCN [2018] 1.137 0.875 129.61 6.801 0.779 0.775
ChatTime [2025] 1.507 1.007 130.75 6.854 0.369 0.435
iTransformer + TaTS (ours) 1.020 0.797 107.11 4.856 0.174 0.218
Comparison with Various Baselines. Besides numerical-only time series models and the MM-TSFLib framework, we compare our TaTS with covariate-based methods N-BEATS [41] and N-HiTS [4], convolution-based TCN [1], and the recent multimodal time series foundation model ChatTime [51]. Although these methods are not specifically designed for time series with paired texts, we adapt them by feeding both time series and text embeddings to covariate-based methods and TCN, while for ChatTime, we concatenate all paired texts into a single long text and perform zero-shot inference. The results are shown in Table 5. The results show that covariate-based and convolution-based models perform comparably or worse than iTransformer, and consistently underperform compared to our TaTS. Although ChatTime achieves competitive results even in a zero-shot setting, highlighting the value of integrating texts when high-quality texts are available, our TaTS outperforms it through supervised learning.
@@ 5.2 Further Analysis
Ablation with Different Text Encoders in TaTS. While GPT-2 was used as the primary text encoder in our main experiments to demonstrate the effectiveness of TaTS, we further evaluate the performance of TaTS with different text encoders, including BERT [7], GPT-2 [43], and LLaMA2 [49]. We utilize the official implementations available on Hugging Face and present the results in Figure 4, with full results provided in Appendix E.7. The results show that TaTS remains robust across different text encoders and consistently outperforms both the uni-modal and MM-TSFLib baselines. Notably, as the size of the language models used in TaTS increases from 110M (BERT) to 1.5B (GPT-2) and further to 7B (LLaMA2), we observe a slight improvement in performance. Investigating the relationship between the text encoder size and TaTS’s effectiveness remains an open direction for future research.
Table 6: Ablation study of TaTS on text-shuffled datasets.
@@
Methods MSE MAE MSE MAE MSE MAE
iTransformer + TaTS + original data 1.020 0.797 107.11 4.856 0.174 0.218
iTransformer + TaTS + corrupted data 1.147 0.831 116.48 5.714 0.208 0.270
iTransformer + uni-modal + original data 1.127 0.843 113.57 5.698 0.203 0.228
Ablation on Corrupted Datasets. To validate that the performance gains stem from useful textual cues, we conducted corruption experiments where the textual information was randomly shuffled across timestamps. All other settings are unchanged, and the results are in Table 6. The results demonstrate that when textual alignment is destroyed, the model’s performance drops significantly to matching or even being worse than the uni-modal baseline. This is expected, as randomly shuffled text acts as noise rather than a meaningful signal.
@@
Due to the inclusion of augmented time series with auxiliary variables from paired texts, the training time per epoch increases slightly, as shown in Figure 5, with average performance of each framework indicated by cross markers. Full results for all datasets are provided in Appendix E.8. However, this marginal efficiency trade-off ($\sim 1\%$ in terms of number of learnable parameters and $\sim 8\%$ in terms of training time) leads to significant improvements ($\sim 14\%$) in forecasting performance, demonstrating the effectiveness of TaTS in enhancing time series modeling with paired texts.
Figure 5: Efficiency comparison on Environment dataset. While TaTS incurs a slight increase in training time due to the augmentation of auxiliary variables, it significantly enhances forecasting accuracy. Full results in Appendix E.8.
@@ 6 Related Work
Numerical-only Time Series Modeling. Recently, various deep learning models have been developed for time series analysis, which can be broadly categorized into three categories. (1) Patch-based models. PatchTST [39] segments time series into subseries-level patches to capture dependencies, while Crossformer [67] employs a two-stage attention mechanism to model both cross-time and cross-variable dependencies efficiently. Autoformer [56] introduces decomposition blocks to separate seasonal and trend-cyclical components. (2) Global representation models. iTransformer [37] utilizes attention over global series representations to capture multivariate correlations. Informer [75] reduces self-attention complexity using ProbSparse self-attention for improved efficiency. Dlinear [63] demonstrates that simple linear regression in the raw space can perform competitively on MTS tasks. (3) Frequency-aware models. FEDformer [77] represents series through randomly selected Fourier components, while FiLM [76] enhances representations with frequency-based layers to reduce noise and accelerate training. In this work, our proposed TaTS is compatible with all of the models listed above.
@@
Time Series with other data sources. Real-world data often coexists and coevolve with auxiliary information [29, 2], such as various modalities [32, 9, 78, 70, 68, 31] and heterogeneous views [73, 69, 28, 30]. In time series analysis, effectively utilizing this auxiliary information remains an emerging and nascent area. In the financial domain, several early works have explored integrating time series with textual data, albeit not in a timestamp-aligned manner, or often leveraging general machine learning models rather than time series-specific architectures. For example, StockNet [60] uses a VAE-like model for chaotic stock-text data, while [45] fuses time series with a single event-related document. BoEC [8] applies a bag-of-economic-concepts approach, and Dandelion [74] leverages multimodal attention for feature aggregation from multiple text sources. Some works also explored time series with vision information [34, 11, 38] for spatio-temporal analysis. Recently, Time-MMD [35] constructs a dataset of time series paired with parallel text, covering multiple domains, which we used in our main experiments.
@@
Large Language Models on Time Series. The rapid advancement of Large Language Models (LLMs) has inspired a new line of research that transforms time series into natural language representations, enabling LLMs to perform downstream tasks [66]. While these approaches demonstrate strong generalization capabilities due to the power of LLMs [12, 3, 16, 61], they also inherit limitations such as hallucination [14] and context length constraints [52, 36]. Notably, a recent work [48] suggests that replacing complex LLM architectures with basic attention layers does not necessarily degrade the performance.
In Appendix B, we also discuss available preprints that are related to this research.
@@ Appendix A Detailed Frequency Analysis Process of Time Series with Paired Texts
Figure 6: Illustration of the frequency analysis process for time series with paired texts in the Economy dataset. Step ①: Compute the average text similarity for each lag $k$. Step ②: Stabilize the time series using first-order differentiation, apply Fourier Transform, and perform Non-Maximum Suppression (NMS) to obtain the amplitude spectrum. Step ③: Visualize the lag similarity of paired texts. Step ④: Stabilize the paired texts, compute the Fourier Transform, and visualize the amplitude spectrum. Step ⑤: Overlay the top-$l$ (here $l$=4) frequencies of paired texts onto the time series amplitude spectrum to highlight shared periodic patterns. All the data transformation operations in this process are periodicity-preserving according to Proposition A.1 and Proposition A.2.
@@
Then, we compute the Fourier Transform [40, 47] of $\Delta\mathbf{X}$ to analyze its frequency components. The Fourier Transform of $\Delta\mathbf{X}$ is defined as:
@@
These top-$l$ frequencies represent the most significant periodic components of the paired texts, revealing their dominant temporal patterns. we overlay the top-$l$ (in the Economy dataset $l$=4) frequencies of the paired texts onto the amplitude spectrum of the time series, as illustrated in Figure 6, subfigure ⑤.

### v3

@@ 5.1 Main Results
Improved Forecasting over Uni-modal Modeling and MM-TSFLib. Table 2 presents the performance results for the time series forecasting task on Time-MMD datasets. For each dataset and time series model, we report the average performance across four different prediction lengths, and the full results for each prediction length are provided in Appendix E. For datasets with relatively few samples, we perform short-term forecasting with prediction lengths of $\{6,8,10,12\}$. In contrast, for datasets with a larger number of samples, we perform long-term forecasting with prediction lengths of $\{48,96,192,336\}$. From the results, compared to uni-modal modeling or MM-TSFLib, our TaTS consistently achieves the best performance across all datasets. Notably, by plugging in TaTS to various existing time series models, it achieves an average performance improvement of over $5\%$ on 6 out of 9 datasets and delivers a remarkable performance boost of over $30\%$ on the largest dataset, Environment. The results also demonstrate that TaTS is highly compatible with a wide range of existing time series forecasting models, consistently delivering performance improvements across all of them in both long-term forecasting and short-term forecasting tasks. We provide a visualization of the performance boost in Appendix E.2, showcasing the improvements achieved by different time series models on each dataset.
Table 3: Imputation task performance. Full results in Appendix E.4. Promotion: percentage reduction by TaTS over the best-performing baseline.
@@
Improved Time Series Imputation. We evaluate the performance of our TaTS framework on the imputation task using the Climate, Economy, and Traffic datasets, each with an imputation length of 24. Though MM-TSFLib only supports forecasting tasks, we extend it to serve as an imputation baseline by applying a similar linear interpolation. We select one representative time series model from each category and present the results in Table 3. The results demonstrate that TaTS consistently enhances the imputation capabilities of existing time series models, achieving improvements of up to $30\%$ compared to baseline methods.
From the above results, effectively leveraging the text modality provides significant benefits when paired texts are available, and our TaTS framework achieves notable improvements over baseline methods that disregard positional information in time-series-paired texts.
Table 4: TaTS compared with several baselines on a variety of datasets. Best results are bolded and second-best results are underlined. Full results in Table 28.
Methods TaTS (ours) TaTS (ours) TaTS (ours) N-BEATS N-HiTS TCN ChatTime GPT4MTS
+ iTransformer + PatchTST + FiLM (2020) (2022) (2018) (2025) (2024)
Datasets
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
Time-MMD: Multimodal Time Series (2024a) Agriculture 0.109 0.229 0.114 0.235 0.135 0.251 3.267 1.458 1.852 1.032 4.168 1.797 0.508 0.447 0.327 0.393
Climate 1.028 0.804 1.004 0.798 0.945 0.772 1.093 0.861 1.103 0.858 1.098 0.866 1.568 1.019 1.127 0.873
Economy 0.008 0.077 0.009 0.079 0.009 0.080 1.010 0.920 0.444 0.584 5.546 2.349 0.049 0.166 0.014 0.096
Energy 0.265 0.376 0.258 0.371 0.271 0.379 0.329 0.424 0.372 0.463 0.430 0.512 0.305 0.417 0.269 0.378
Environment 0.267 0.369 0.273 0.371 0.272 0.371 0.518 0.573 0.522 0.583 0.854 0.713 0.580 0.594 0.348 0.422
Health 1.315 0.744 1.283 0.753 1.421 0.838 1.660 0.938 1.666 0.898 1.938 0.970 1.668 0.909 1.766 0.875
Security 112.054 5.151 109.693 5.019 107.850 4.736 130.065 6.618 138.124 6.978 136.596 6.873 133.106 6.887 119.407 5.198
Social Good 0.987 0.452 0.972 0.465 1.104 0.626 1.316 0.665 1.272 0.649 1.314 0.961 1.264 0.652 1.414 0.559
Traffic 0.187 0.217 0.172 0.209 0.176 0.248 0.347 0.464 0.268 0.382 0.708 0.744 0.363 0.429 0.195 0.246
FNSPID: Company Stock Price (2024) Delta Airlines (DAL) 0.087 0.197 0.086 0.192 0.095 0.201 0.286 0.444 0.226 0.379 0.294 0.466 0.098 0.202 0.093 0.201
IBM (IBM) 0.564 0.501 0.550 0.490 0.891 0.695 1.105 0.777 1.215 0.806 1.936 1.123 0.602 0.536 0.639 0.525
JPMorgan (JPM) 1.693 0.970 1.872 0.990 2.513 1.096 2.419 1.175 3.426 1.232 3.764 1.711 2.037 1.043 2.133 1.122
NVIDIA (NVDA) 0.043 0.141 0.048 0.156 0.050 0.174 0.272 0.434 0.122 0.262 0.457 0.574 0.053 0.161 0.054 0.159
Pfizer (PFE) 0.326 0.416 0.347 0.422 0.448 0.477 0.676 0.572 0.824 0.663 0.628 0.559 0.408 0.477 0.369 0.439
Tesla (TSLA) 0.142 0.281 0.158 0.297 0.110 0.244 0.188 0.332 0.254 0.416 3.476 1.851 0.158 0.300 0.181 0.303
FNF: Forecast with News (2024b) Bitcoin Price 2.609 1.112 2.339 1.045 2.775 1.142 141.738 10.355 159.019 10.535 57.826 6.744 3.590 1.344 2.721 1.152
Web Traffic 17.744 2.712 17.964 2.653 18.110 2.772 22.233 3.017 23.263 3.011 21.626 2.842 20.748 2.909 19.260 2.827
Electricity Demand 0.280 0.396 0.254 0.358 0.266 0.377 0.416 0.496 0.364 0.468 0.516 0.564 0.532 0.566 0.296 0.395
Table 5: TaTS Improvements positively correlated with TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets. ”Original-Shuffled $T$ Ratio” means the ratio of TT-Wasserstein on the original dataset to TT-Wasserstein on the shuffled dataset. ”TaTS Improvement” means the average improvement of TaTS over uni-modal time series modeling.
@@
Original-Shuffled TT Ratio (%) 26.8 72.4 22.3 91.5 38.2 34.6 97.1 85.6 83.6
TaTS Improvement (%) 20.70 18.03 64.80 4.05 10.99 16.78 3.62 19.51 36.00
Comparison beyond Uni-modal Modeling and MM-TSFLib. Besides numerical-only time series models and the MM-TSFLib framework, we compare our TaTS with several other baselines on three dataset sources, including Time-MMD (Liu et al., 2024a), FNSPID (Dong et al., 2024), and FNF (Wang et al., 2024b). Although N-BEATS, N-HiTS and TCN are not specifically designed for time series with paired texts, we adapt them by feeding both time series and text embeddings to them. We also include several baselines that explicitly incorporate multimodal information.. For ChatTime, we concatenate all paired texts into a single long text and perform zero-shot inference. For GPT4MTS, we convert the datasets we used into their prescribed input format and directly apply their pipeline. The results averaged from multiple prediction lengths are shown in Table 4, with full results in Table 28. The results show that covariate-based and convolution-based models perform comparably or worse than iTransformer, and consistently underperform compared to our TaTS. Although ChatTime achieves competitive results even in a zero-shot setting, highlighting the value of integrating texts when high-quality texts are available, our TaTS outperforms it through supervised learning. We further compare TaTS with several multimodal methods from concurrent preprints that also evaluate on Time-MMD. Using their reported results (Table 12), we find that TaTS elevates standard time-series backbones into highly competitive models, rivaling recent approaches with substantially more complex designs.
@@ 5.2 Further Analysis
Ablation with Different Text Encoders in TaTS. While GPT-2 was used as the primary text encoder in our main experiments to demonstrate the effectiveness of TaTS, we further evaluate the performance of TaTS with different text encoders, including BERT (Devlin et al., 2019), and LLaMA2 (Touvron et al., 2023). We utilize the official implementations available on Hugging Face and present the results in Figure 4, with full results provided in Appendix E.7. The results show that TaTS remains robust across different text encoders and consistently outperforms both the uni-modal and MM-TSFLib baselines. Notably, as the size of the language models used in TaTS increases from 110M (BERT) to 1.5B (GPT-2) and further to 7B (LLaMA2), we observe a slight improvement in performance. Investigating the relationship between the text encoder size and TaTS’s effectiveness remains an open direction for future research.
Table 6: TaTS Ablation on Alternative Modality Fusion Architectures.
Dataset Climate Security Traffic # Parameters
TaTS Settings MSE MAE MSE MAE MSE MAE Fuser Total
MLP projection (original) 0.992 0.791 109.865 4.968 0.178 0.224 74988 6396658
gated residual 0.998 0.794 110.254 4.965 0.177 0.227 9396 6331066
cross-attention 0.989 0.792 109.729 4.976 0.178 0.221 99340 6421010
Ablation with Other Design Choices to Combine Modalities. In our main experiments, TaTS transforms the paired texts into auxiliary variables through embedding, projection and concatenation. Our TaTS framework is readily extensible to other fusion design choices. We evaluate two alternative architectures that replace the MLP projection with: (a) a gated residual and (b) a cross-attention module between the time series and text embeddings. We report the average MSE and MAE across TaTS with iTransformer, PatchTST, and FiLM in Table 6. From the results, these alternative mechanisms achieve similar performance to the MLP-based design. One possible explanation is that linear projections are already highly competitive for time-series representation learning, especially when paired with strong backbone forecasters that account for the majority of the model parameters. We leave the exploration of more fine-grained multimodal fusion designs as an interesting direction for future work.
Table 7: TaTS Ablation on text-shuffled datasets.
@@
Settings MSE MAE MSE MAE MSE MAE
iTransformer + TaTS + original data 1.028 0.804 112.05 5.151 0.187 0.217
iTransformer + TaTS + corrupted data 1.242 0.895 117.82 5.767 0.223 0.265
iTransformer + uni-modal + original data 1.183 0.871 115.94 5.660 0.213 0.238
Ablation on Corrupted Datasets. To validate that the performance gains stem from useful textual cues, we conduct corruption experiments where the textual information was randomly shuffled across timestamps. All other settings are unchanged, and the results are in Table 7. When textual alignment is destroyed, the performance drops significantly to matching or even being worse than the uni-modal baseline. This is expected, as randomly shuffled text acts as noise rather than a meaningful signal. Yet our TaTS stays relatively robust because it can learn to assign a small weight to textual information during optimization as a mitigation to noisy texts.
Table 8: TaTS Ablation on text-dropped datasets.
Dataset Climate Security Traffic
Settings MSE MAE MSE MAE MSE MAE
TaTS + original data 1.028 0.804 112.05 5.151 0.187 0.217
TaTS + 10% text randomly dropped 1.032 0.809 112.94 5.372 0.194 0.239
TaTS + 25% text randomly dropped 1.056 0.818 114.59 5.231 0.203 0.254
Ablation on Text-Randomly-Dropped Datasets. Similar to corruption ablations, we conduct experiments with randomly dropped texts. Dropped texts are filled with “no information available.” The results are in Table 8. While performance degrades slightly as more text is randomly dropped, TaTS remains effective, performing comparably to MM-TSF Lib even with 25% of the text missing. This experiment shows that whenever textual information is available, our TaTS could benefit from it to improve time series modeling.
Table 9: Dropping extremely noisy texts mitigates negative effects from them.
Dataset Climate Security Traffic
Settings MSE MAE MSE MAE MSE MAE
TaTS + corrupted data 1.242 0.895 117.82 5.767 0.223 0.265
TaTS + 60% corrupted data 1.218 0.879 117.43 5.730 0.215 0.260
TaTS + 20% corrupted data 1.201 0.878 116.68 5.713 0.216 0.255
uni-modal + original data 1.183 0.871 115.94 5.660 0.213 0.238
Mitigating Negative Effects of Extremely Noisy Texts. Our ablation on corrupted datasets shows that highly noisy text can cause TaTS to even slightly underperform numerical-only modeling. To address this, we evaluate a simple mitigation strategy: randomly dropping a portion of the noisy texts and replacing them with “no information available.” We test drop rates of 40% and 80%. As shown in Table 9, dropping noisy texts allows TaTS to recover performance close to the unimodal baseline, demonstrating the effectiveness of this strategy in handling unreliable text. Practitioners can refer to Appendix B.2 for how TaTS can adapt to scenarios with low-quality texts.
@@
Due to the inclusion of augmented time series with auxiliary variables from paired texts, the training time per epoch increases slightly, as shown in Figure 5, with average performance of each framework indicated by cross markers. Full results for all datasets are provided in Appendix E.8. Notably, this marginal efficiency trade-off ($\sim 1\%$ in terms of number of learnable parameters and $\sim 8\%$ in terms of training time) leads to significant improvements ($\sim 14\%$) in forecasting performance.
Figure 5: While TaTS incurs a slight increase in training time due to augmented auxiliary variables, it significantly improves forecasting. Full results in Appendix E.8.
@@ 6 Related Work
Numerical-only Time Series Modeling. Recently, various deep learning models have been developed for time series analysis, which can be broadly categorized into three categories. (1) Patch-based models. PatchTST (Nie et al., 2023) segments time series into subseries-level patches to capture dependencies, while Crossformer (Zhang and Yan, 2023) employs a two-stage attention mechanism to model both cross-time and cross-variable dependencies efficiently. Autoformer (Wu et al., 2021) introduces decomposition blocks to separate seasonal and trend-cyclical components. (2) Global representation models. iTransformer (Liu et al., 2024c) utilizes attention over global series representations to capture multivariate correlations. Informer (Zhou et al., 2021) reduces self-attention complexity using ProbSparse self-attention for improved efficiency. Dlinear (Zeng et al., 2023) demonstrates that simple linear regression in the raw space can perform competitively on MTS tasks. (3) Frequency-aware models. FEDformer (Zhou et al., 2022b) represents series through randomly selected Fourier components, while FiLM (Zhou et al., 2022a) enhances representations with frequency-based layers to reduce noise and accelerate training. In this work, our proposed TaTS is compatible with all of the models listed above.
@@
Time Series with other data sources. In the financial domain, several early works have explored integrating time series with textual data, albeit not in a timestamp-aligned manner, or often leveraging general machine learning models rather than time series-specific architectures. For example, StockNet (Xu and Cohen, 2018) uses a VAE-like model for chaotic stock-text data, while (Rodrigues et al., 2019) fuses time series with a single event-related document. BoEC (Farimani et al., 2021) applies a bag-of-economic-concepts approach, and Dandelion (Zhou et al., 2020) leverages multimodal attention for feature aggregation from multiple text sources. Some works also explored time series with vision information (Liu and Cai, 2012; Gerard et al., 2023; Lütjens et al., 2024) for spatio-temporal analysis. Recently, Time-MMD (Liu et al., 2024a) constructs a dataset of time series paired with parallel text, covering multiple domains, which we used in our main experiments.
@@
Large Language Models on Time Series. The rapid advancement of Large Language Models (LLMs) has inspired a new line of research that transforms time series into natural language representations, enabling LLMs to perform downstream tasks (Zhang et al., 2024c). While these approaches demonstrate strong generalization capabilities due to the power of LLMs (Gruver et al., 2023; Cao et al., 2024; Jin et al., 2024; Xue and Salim, 2024), they also inherit limitations such as hallucination (Huang et al., 2023) and context length constraints (Wang et al., 2024a; Liu et al., 2024b). Notably, a recent work (Tan et al., 2024) suggests that replacing complex LLM architectures with basic attention layers does not necessarily degrade the performance. Another emerging line of work focuses on generating texts from time series, with applications in temporal reasoning and time-series question answering (Chang et al., 2025). MTBench (Chen et al., 2025) provides a comprehensive evaluation framework for assessing whether LLMs can jointly reason over structured numerical trends and unstructured textual narratives. TSAIA (Ye et al., 2025) benchmarks LLMs’ multi-step temporal reasoning capabilities. TimeXL (Jiang et al., 2025) further introduces collaborating LLM agents to generate interpretable natural-language explanations alongside time-series forecasts. One notable future direction is to enable latent communications (Zou et al., 2025a; b) between the time series model and language model to build general time series intelligence.
@@ Appendix A Detailed Frequency Analysis Process of Time Series with Paired Texts
Figure 6: Illustration of the frequency analysis process for time series with paired texts in the Economy dataset. Step ①: Compute the average text similarity for each lag $k$. Step ②: Stabilize the time series using first-order differentiation, apply Fourier Transform, and perform Non-Maximum Suppression (NMS) to obtain the amplitude spectrum. Step ③: Visualize the lag similarity of paired texts. Step ④: Stabilize the paired texts, compute the Fourier Transform, and visualize the amplitude spectrum. Step ⑤: Overlay the top-$l$ (here ll=4) frequencies of paired texts onto the time series amplitude spectrum to highlight shared periodic patterns. All the data transformation operations in this process are periodicity-preserving according to Proposition A.1 and Proposition A.2.
@@
Then, we compute the Fourier Transform (Nussbaumer, 1982; Sneddon, 1995) of $\Delta\mathbf{X}$ to analyze its frequency components. The Fourier Transform of $\Delta\mathbf{X}$ is defined as:
@@
These top-$l$ frequencies represent the most significant periodic components of the paired texts, revealing their dominant temporal patterns. we overlay the top-$l$ (in the Economy dataset ll=4) frequencies of the paired texts onto the amplitude spectrum of the time series, as illustrated in Figure 6, subfigure ⑤.
@@
A.1 Estimation Stability of TT-Wasserstein
As TT-Wasserstein is an empirically estimated statistical metric, we provide a detailed sensitivity and robustness analysis in this appendix. Our default configuration uses a rectangular window, $\ell_{1}$-normalization of spectral amplitudes, and a frequency grid determined by the full length of the time series (native FFT resolution). To assess the stability of TT-Wasserstein, we perform ablations that vary each of these components: windowing schemes, normalization strategies, and frequency resolutions, while keeping all other settings fixed. The results are summarized in Table 10, where each row reflects the effect of a single deviation from the default configuration. From the results, our TT-Wasserstein is stable across various configuration choices. We also implemented block bootstrap with a bootstrap sample size of 10 for time series and texts to compute the confidence intervals for the default settings in Table 11.
Table 10: TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets. Our TT-Wasserstein is generally stable on different frequency computation configurations.
Dataset Monthly Sampled Weekly Sampled Daily Sampled
Default 0.026 0.025 0.022 0.049 0.027 0.035 0.307 0.233 0.302
Hann Window 0.031 0.029 0.016 0.041 0.040 0.057 0.304 0.243 0.306
Hamming Window 0.054 0.025 0.016 0.041 0.037 0.046 0.305 0.242 0.305
Blackman Window 0.061 0.029 0.020 0.036 0.043 0.055 0.297 0.245 0.298
Minmax Normalization 0.032 0.035 0.025 0.043 0.031 0.049 0.345 0.291 0.332
Log Normalization 0.016 0.060 0.046 0.049 0.050 0.048 0.267 0.170 0.286
Zero-padded frequency resolution ($2\times$) 0.046 0.058 0.052 0.051 0.028 0.050 0.280 0.187 0.292
Down-sampled frequency resolution ($\frac{1}{2}\times$) 0.011 0.026 0.039 0.050 0.051 0.025 0.259 0.197 0.298
Table 11: Standard deviation of default TT-Wasserstein measure of Time-MMD (Liu et al., 2024a) datasets by block bootstrap.
Dataset Monthly Sampled Weekly Sampled Daily Sampled
Default $\pm$0.06 $\pm$0.04 $\pm$0.03 $\pm$0.03 $\pm$0.04 $\pm$0.06 $\pm$0.21 $\pm$0.16 $\pm$0.12